OpenStudy (anonymous):

Assume that y varies inversely with x. If y = -3 when x = 2.4, what is y when x = 2?

OpenStudy (psymon):

Okay, so inversely means we have the set up of \[y = \frac{ k }{ x }\]Your first job is to always use y and x to help you find k, then usually you use k and x to find y in the second portion. SO based on how the inverse format should look, I an plug in -3 for y and 2.4 for x and solve for k \[-3 = \frac{ k }{ 2.4 }\] Multiply both sides by 2.4 \[-3*2.4 = k \implies k = -7.2\]Now we got k, so we can do the second part of the problem. We use x = 2 and k = -7.2 and solve for y, using the same format as before. \[y = \frac{ k }{ x } \implies y = \frac{ -7.2 }{ 2 } \implies y = -3.6\]

OpenStudy (anonymous):

ily

OpenStudy (psymon):

Lol, np then I guess xD